OpenStudy (anonymous):

what is (sqrt)605w24

OpenStudy (anonymous):

Can you type using equation editor?

OpenStudy (anonymous):

\[\sqrt{605w ^{24}}\]

OpenStudy (anonymous):

\[\sqrt{605}*w^{12}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

605=121*5 \[11\sqrt{5}*w^{12}\]

OpenStudy (anonymous):

thanks